OpenStudy (anonymous):

(secx-cosx)/tanx I get that it becomes ((1/cosx)-(cosx))/(sinx/cosx), but how does it become 1-cos^2x/cosx?

OpenStudy (amistre64):

\[\frac1c-c=?\]

OpenStudy (amistre64):

looks to be a sin denominator in the end, nota cos ... but i cant see it that clearly yet

OpenStudy (amistre64):

\[(\frac1c-c)\frac cs=\frac{1-c^2}{s}=\frac{s^2}{s}=s\]

OpenStudy (anonymous):

in that case, where did the cosine denominator in the (1/c-c) go? I get that you would multiply by the reciprocal of tangent, or cotangent, but where did the cosine go?